OpenStudy (anonymous):

Solve the equation: 4y^2-4y-19=0

OpenStudy (anonymous):

solve dis question pls If A : B = 3 : 4, C : B = 5 : 4, C : D = 10 : 9, then A : B : C : D is

OpenStudy (campbell_st):

needs the general quadratic formula... I'd expect.

OpenStudy (anonymous):

The answer is suppose to be: y=[1(+or-)2√5]/2, but I have: y=[-4(+or-)8√5]/8. Can someone please explain what I'm doing wrong?

OpenStudy (campbell_st):

well its \[x = \frac{4 \pm \sqrt{4^2 - 4 \times 4 \times (-19)}}{2 \times 4} = \frac{ 4 \pm \sqrt{320}}{8}\] just simplify

OpenStudy (campbell_st):

then \[x = \frac{4 \pm 8\sqrt{5}}{8} = \frac{1 \pm 2\sqrt{5}}{2}\]

OpenStudy (campbell_st):

remove a common factor of 4 from every term.

OpenStudy (anonymous):

What I did wrong was plug the 4 in as a negative. Thank you!